OpenStudy (anonymous):

whats the mean and the range 20 14 29 39 33 25 19 35

OpenStudy (anonymous):

mean = average ( add them all and divide by the total number ) range = highest data - lowest data

OpenStudy (anonymous):

mean = (20 + 14 +29 + 39 + 33 + 25 + 19 + 35 ) / 8 = 26.75 range = (39-14) = 25

OpenStudy (anonymous):

EDIT : mean = average ( add them all and divide by the total number of data values) thought I should specify it a bit better, because "dividing by total number" isnt very good explaination